### v1

# Making FETCH! Happen: Finding Emergent Dog Whistles Through Common Habitats

Abstract
WARNING: This paper contains content that maybe upsetting or offensive to some readers.
Dog whistles are coded expressions with dual meanings: one intended for the general public (outgroup) and another that conveys a specific message to an intended audience (ingroup). Often, these expressions are used to convey controversial political opinions while maintaining plausible deniability and slip by content moderation filters. Identification of dog whistles relies on curated lexicons, which have trouble keeping up to date. We introduce FETCH!, a task for finding novel dog whistles in massive social media corpora. We find that state-of-the-art systems fail to achieve meaningful results across three distinct social media case studies. We present EarShot, a novel system that combines the strengths of vector databases and Large Language Models (LLMs) to efficiently and effectively identify new dog whistles. [^1] Code for analysis and evaluation can be found at https://github.com/KuleenS/FETCH-Dog-Whistle
1 Introduction
white house petition: end dual citizens from serving in us and state governments
abcAnonymous Gab User
When a person writes dual citizens it may seem like an innocuous phrase, but many understand this as an antisemitic reference. The use of the phrase builds on centuries old accusations that question the loyalty of Jewish people towards state governments, which in recent decades has become a claim that all American Jews are double citizens of Israel and the United States.
This phrase is a dog whistle, which is defined as an “expression that send one message to an outgroup while at the same time sending a second (often taboo, controversial, or inflammatory) message to an ingroup Henderson and McCready (2018).” Dog whistles are a common mainstay in today’s turbulent political climate. For example, they prominently featured in the 2016 United States Presidential Election both on social media and in person Tilley et al. (2020); Drakulich et al. (2020).
Figure 1: Comparison between dog whistle detection and our method of dog whistle discovery
As dog whistles are a common way to transmit hate speech to a wide audience both digitally and offline, researchers have started to develop methods to detect them. However, dog whistles are notoriously difficult to find as they often slip by content moderation, toxicity, and hate speech filters at a high rate because of their benign second meaning Mendelsohn et al. (2023). Dual citizenship isn’t an issue, but “dual citizen” in the above context has a specific meaning.
Currently, detecting dog whistles relies on manually curated lexicons, which can be effective but are labor-intensive to create and maintain. They cannot adapt to the dynamic evolution of language. For example (Figure 1), current methods can distinguish between the two meanings of the word “cosmopolitan”: its use as the popular mixed drink and its use as a dog whistle to denigrate a multicultural and intellectual society. However, they miss a hypothetical linguistic shift of the term “cosmopolitan” to the shortening of “cosmos.”
We consider the task of dog whistle discovery, a novel task where systems identify emerging or previously unknown dog whistles. To support this effort, we create FETCH! or Finding Emergent Dog Whistles Through Common Habitats, a benchmark designed to evaluate models’ ability to detect novel dog whistles across diverse settings. Using the FETCH! benchmark, we evaluate three state of the art methods across three different social media case studies. Our evaluation reveals that these methods perform poorly, achieving an F-score of less than 5% across all case studies.
We introduce our own method: EarShot. This method finds new dog whistles by leveraging the power of sentence embeddings and Large Language Models (LLMs). EarShot demonstrates significant improvements, with F-score increases ranging from $2$ to $20$ points across the social media case studies. We make the following key contributions:
  • Development of FETCH! benchmark for dog whistle discovery;
  • Rigorous evaluation of multiple state-of-the-art methods on FETCH!;
  • Creation of EarShot, a novel architecture achieving SoTA performance on FETCH!.
2 Related Work
2.1 Implicit Hate Speech Detection
Detecting novel dog whistles can be thought of as a form of hate speech detection, specifically implicit hate speech detection. Implicit hate speech can be defined as hate speech that uses coded or indirect language Waseem et al. (2017). Many studies have focused on this task through creating new datasets Hartvigsen et al. (2022); ElSherief et al. (2021); Ocampo et al. (2023); Breitfeller et al. (2019), novel developing detection methods Zhang et al. (2024); Gao et al. (2017), and methods to explain the hate speech Qian et al. (2019). However, these methods focus on the overall text or post rather than the specific words and phrases.
2.2 Dog Whistles
Other fields like linguistics, political science, and other social sciences have researched the use of dog whistles and their appearance through various different forms of analysis. Linguistic methods have focused on the semantics and pragmatics of their use Henderson and McCready (2018); Torices (2021); Khoo (2017); Saul (2018); Bhat and Klein (2020); Quaranto (2022) and further extended their work using agent-based simulations of the emergence of new dog whistles Henderson and McCready (2020). Historians have focused on showing the evolution and propagation of racial and dog whistle speech for political purposes and their lasting effects on society Mendelberg (2001); Haney-López (2014). Sociologists and psychologists have focused on surveying attitudes of people before and after reading messages with implicit racial appeals like dog whistles Wetts and Willer (2019); Albertson (2015). Other sociologic work has studied the development and changes of dog whistles over time on the web Åkerlund (2022). Political science has investigated the use of dog whistles and how they change public opinions and policies surrounding controversial topics like sanctuary cities and other criminal justice policies Lasch (2016); Hurwitz and Peffley (2005); Goodin and Saward (2005).
2.3 Euphemism Detection
Euphemism detection focuses on finding novel euphemisms, terms or phrases used to substitute more offensive terms in order to downplay its unpleasantness, a semantic phenomenia similar to dog whistles. Typically, the task is structured as given a set of seed euphemisms, the goal of the task is to find other euphemisms in a provided corpora, a process similar to lexicon induction Lee et al. (2022).
Early works in euphemism detection often focused on using Word2Vec Mikolov et al. (2013) to find euphemisms. In one work, the authors trained a Word2Vec model on the corpus of interest, and then, they used the similarities between the word vectors to obtain similar words to their seed words Magu and Luo (2018). Other works further improved on this method combining both Word2Vec and dependency based word embeddings Levy and Goldberg (2014); Taylor et al. (2017), using two different Word2Vec models on two different datasets and finding the words with the greatest difference in embeddings HADA et al. (2020); Yuan et al. (2018); Zhao et al. (2016), or using the difference between the two Word2Vec learned matrices Aoki et al. (2017). Other works used methods combining Word2Vec methods with additional user and search behaviors of users Yang et al. (2017). Later works further refined these methods by using both more powerful Masked Language Models like BERT Devlin et al. (2019); Zhu et al. (2021) and SpanBERT Joshi et al. (2020); Zhu and Bhat (2021) both in a zeroshot setting Maimaitituoheti et al. (2022) and a supervised setting and Large Language Models like GPT-3 Brown et al. (2020) in a zeroshot setting Keh (2022).
2.4 Dog Whistle Detection
Relatively few studies have addressed the detection of dog whistles. Among these, three notable works focus specifically on Swedish dog whistles, employing both embedding-based methods and analyzing semantic shifts Hertzberg et al. (2022); Boholm and Sayeed (2023); Boholm et al. (2024). In English, only two studies have focused on detecting dog whistles. In Mendelsohn et al. (2023), they curated set of English dog whistles to assess both large language models’ (LLMs) ability to surface these terms and the impact on hate speech detectors. However, they only do a small analysis with a handful of examples to see if LLMs can detect novel dog whistles. In Kruk et al. (2024), they introduced the largest dataset of posts containing dog whistles, filtered from sources such as Reddit and the Congressional Record by LLMs. In addition to the creation of the dataset, they conducted a preliminary test of LLMs’ capabilities in detecting novel dog whistles.
3 Task Description
We propose a novel task: Finding Emergent Dog Whistles in Common Habitats or FETCH!. For the task, a system is provided with a corpus and a set of initial seed dog whistles. The corpus might be social media posts, forums, website pages, etc. The initial set of seed dog whistles is a list of known dog whistles that occur within the corpus. The goal of the task is to use the initial seed words and the corpus to discover other dog whistles.
3.1 Case Studies
We describe the three example case studies that represent real-world or synthetic scenarios for our task.
3.1.1 Synthetic (Reddit)
Our first scenario is an idealized setting where every post contains an associated dog whistle. We use the Silent Signals dataset Kruk et al. (2024), which includes 16,000 posts scraped from Reddit. Each post in the dataset contains a dog whistle identified by GPT-4 OpenAI et al. (2024).
3.1.2 Balanced (Gab)
The balanced scenario represents a context with a higher-than-average prevalence of dog whistles with large amount of posts. For this scenario, we chose to use data from Gab, an alt-right alternative to Twitter where dog whistles and other forms of hate speech are more frequent. The data was collected over a short period using a web scraper that sequentially downloaded post IDs. This process yielded approximately 300,000 posts.
3.1.3 Realistic (Twitter)
The realistic scenario reflects a typical online environment where dog whistles are sparse, and most posts are benign. For this scenario, we selected X (formerly Twitter) as the platform, as it contains some hate speech but removes much of it through content moderation. Data were collected using the $1$% public Twitter API stream between January 2017 and December 2022. Given the dataset’s massive size (approximately $1$ billion tweets), we subsampled it to make it manageable. We randomly sampled 6 million English language tweets without dog whistles and $1$ million tweets containing dog whistles identified using regular expression from a set of ground truth dog whistles. This process resulted in a dataset of 7 million tweets.
3.2 Seed Dog Whistles
We use the dog whistle dataset released by Mendelsohn et al. (2023) as our ground truth. The dataset is a hand-curated taxonomy of around 340 root dog whistles from community wikis and academic sources. Each root dog whistle comes with the usage of the word, the group it is targeting, the group that often uses it, definitions, and other surface forms of the root dog whistle. We use these dog whistles for the Realistic and Balanced scenarios as the Synthetic scenario has its own set of dog whistles. Table 1 lists a summary of the number of dog whistles found in each scenario. The full list appears in Appendix A.
Scenario Balanced Synthetic Realistic
Total Dog Whistles Found 77 298 177
Reference Dog Whistles Total 340 298 340
Table 1: Number of dog whistles root forms found in each scenario
For each corpora, we divide the dog whistles into train/test splits (20% / 80%), creating a set of seed dog whistles and dog whistles to find. We split the dog whistles using stratified sampling based on how the length of phrase in $n$-grams. We do a stratified split so as not to unfairly punish methods that can identify larger $n$-grams. After splitting, we add the surface forms as well to each split in order to cover all variants of the dog whistles.
3.3 Metrics for Evaluation
We use precision, recall and F-score to measure the performance of different methods for our task. Precision is compute as normal, but we modify Recall to normalize by only the positive terms that exist in the corpus (Data Potential Recall, DPR). Due to the multiple surface forms of dog whistles provided in the dataset, all metrics count a positive match for a single root dog whistle as returning an exact match of one of the surface forms or the root itself.
We report $F_{0.5}$ scores using DPR and Precision instead of $F_{1}$ to preference methods with a high precision to reduce the load on a human reviewer.
4 Experimental Setup
4.1 Models
We test four different methods on FETCH!: Word2Vec/Phrase2Vec Mikolov et al. (2013), MLM Zhu et al. (2021), EPD Zhu and Bhat (2021), and our method: EarShot.
4.1.1 Word2Vec/Phrase2Vec
Word2Vec and Phrase2Vec are widely used models for euphemism detection due to their lightweight nature and fast computation. These models capture semantic relationships between words or phrases, enabling the detection of euphemistic language. In our implementation of these models for dog whistle discovery, we adopt a methodology similar to Magu and Luo (2018).
To create the training data for the Word2Vec/Phrase2Vec Mikolov et al. (2013), we preprocessed the text to remove all stop words and lemmatize the words using littlebird DeLucia (2020). After this preprocessing step, we replaced URLs, hashtags, retweets, and mentions with special symbols like @USER and HTTPURL. We also replaced emojis with their official English aliases using the emoji package Kim and Wurster (2024). Finally, we tokenized the result with the NLTK Tweet Tokenizer Bird and Loper (2004).
To construct the Phrase2Vec models, we utilized the Gensim Phraser module Rehurek and Sojka (2011) to merge common $n$-gram phrases in the original dataset, generating $2$-gram and $3$-gram versions of the models. We only train up to $3$-gram models as a majority of dog whistles in our dataset are $3$-grams and smaller. We then trained models using Gensim with a maximum vocabulary size of 500,000 and a context window of 5 with a vector dimension of 100 for 10 epochs
To generate the predictions, we provided the Word2Vec/Phrase2Vec model the list of seed dog whistles and retrieved the top 10 most similar words to each seed word. Then, we queried for the top 10 most similar words to those words. We continued this process 10 times creating 10 levels of predictions.
4.1.2 MLM
Euphemism detection relies heavily on context, but static embeddings from Word2Vec/Phrase2Vec fail to capture this. To address this, Zhu et al. (2021) introduces a model leveraging contextual embeddings from Masked Language Models like BERT. The MLM system first samples 2000 sentences from the input dataset containing the seed dog whistles. We sampled 2000 sentences as it is default parameter provided. For each sampled sentence, it masks out the seed dog whistle in that sentence and uses a Masked Language Model to fill in the mask by creating a distribution over possible replacements and then returning the top-$k$ most likely. We test this method using Bernice DeLucia et al. (2022) and BERTweet Nguyen et al. (2020), two MLMs that have already been adapted to social media data. We test the method at thresholds of $\{50,100,200,400,800,1600,3200,6400,12800,\\ 25600\}$.
Figure 2: Flowchart for the EarShot System. Two paths can be taken after doing the nearest neighbor lookup. Path (1) is the more computationally expensive system that asks directly prompts a LLM for an answer. Path (2) is the cheaper method that leverages keyword extraction models and binary prediction.
4.1.3 EPD
The Euphemistic Phrase Detector (EPD) Zhu and Bhat (2021) builds upon the MLM model, addressing its limitation of restricting predictions to a single token from the tokenizer. To solve this issue, the EPD model generates possible phrases using Autophrase Shang et al. (2017), a statistical phrase extractor to choose important phrases from a corpus. Next, it trains a unigram Word2Vec model Mikolov et al. (2013) on the same corpus. It uses this model to filter down the phrases by selecting the top 1000 most similar phrases from AutoPhrase to the seed dog whistles. Finally, similar to MLM method, it samples 2000 sentences from the input dataset that contain the seed dog whistles. For each sampled sentence, it replaces the seed dog whistle with each phrase kept from the previous stage and uses a Masked Language Model to calculate the probability of the sentence with the replaced phrase. They then sort the sentences via their probability and select the top-$k$ sentences and return the corresponding inserted phrase as predictions. We test this method using SpanBERT Joshi et al. (2020) as our Masked Language Model. We test the method at increasing thresholds of $\{50,100,200,400,800,1600,3200,6400,12800,\\ 25600\}$.
4.1.4 EarShot
We propose a new method called EarShot. EarShot consists of three stages.
In the first stage, all posts in the corpora are turned into vectors using a sentence transformer Reimers and Gurevych (2019) and put into a vector database. We opt to use the all-MiniLM-L6-v2 as our sentence transformer as it is one of the fastest models with high quality embeddings. We use ChromaDB Chroma (2024) as our vector database.
In the second stage, we obtain the vectors for the posts containing the seed dog whistles and find the closest vector to each post vector that are not the post themselves. By identifying closet neighbors, we aim to capture posts that are likely related in meaning, sentiment, or intent, even if they do not share exact word matches.
After this step, there are two different pipelines we try: DIRECT and PREDICT, labeled (1) and (2) in Figure 2 respectfully.
For the DIRECT pipeline, we pass all the posts to a LLM to directly extract the dog whistles from the posts provided from the vector lookup. Similarly to Kruk et al. (2024), the model is prompted to list all dog whistles in a structured JSON format. The prompt is shown in Table 7 in Appendix B. We evaluate three LLMs: meta-llama/Meta-Llama-3.1-8B-Instruct (LLaMa 8B) Dubey et al. (2024), mistralai/Mistral-7B-Instruct-v0.3 (Mistral 7B) Jiang et al. (2023), and meta-llama/Llama-2-13b-chat-hf (LLaMa 13B) Touvron et al. (2023). We choose to test LLaMa and Mistral as they some of the best performing open source LLMs.
For the PREDICT pipeline, we have two stages. In the first stage, we have a filtering step. We filter the posts using either a BERT model or a LLM. For the BERT based filtering, we use an off the shelf hate speech or toxicity detection model to classify the posts. We remove all the posts that are labeled not toxic or hate speech according to the model as dog whistles are commonly found in hate speech. We call this method filtering BERT-PREDICT. We test three popular hate speech and toxicity classifiers: tomh/toxigen_hatebert (ToxiGen BERT) Hartvigsen et al. (2022), facebook/robertahate-speech-dynabench-r4-target (RoBERTa R4) Vidgen et al. (2021), and Hate-speechCNERG/bert-base-uncased-hatexplain (HateXplain BERT)Mathew et al. (2022).
For the LLM based filtering, we prompt the model to reply yes or no if there is a dog whistle in the post. We structure our prompt similarly to Silent Signals Kruk et al. (2024), shown in Table 7 in Appendix B. We keep all the posts that the model explicitly says yes to using a substring search. We call this filtering method LLM-PREDICT. We test the same three models used in the DIRECT pipeline.
After the filtering step, we pass the posts to a keyword extractors and return the top-$k$ predictions of them. We use five different popular keyword extraction algorithms: KeyBERT Grootendorst (2020), RAKE Rose et al. (2010), YAKE, Campos et al. (2018), TextRank Mihalcea and Tarau (2004), and TF-IDF. If the keyword extractor requires a specific range of $n$-grams to return, we test ranges of $1$-gram, 11112222-grams, and 11113333-grams. We report the performance of the method at the same thresholds in MLM and EPD.
5 Results
5.1 Word2Vec/Phrase2Vec
Scenario Model Level Prec DPR $F_{0.5}$
Balanced Unigram 1 1.64 4.12 1.86
Bigram 1 0.83 6.17 1.00
Trigram 1 0.83 6.17 1.00
Synthetic Unigram 1 5.50 8.40 5.91
Bigram 2 2.82 21.85 3.42
Trigram 1 3.24 7.98 3.67
Realistic Unigram 1 2.97 1.47 2.47
Bigram 1 0.89 1.47 0.97
Trigram 1 1.03 1.47 1.10
Table 2: Results for best Word2Vec/Phrase2Vec models by $F_{0.5}$. Prec is Precision. DPR is Data Potential Recall. Best scores for across each dataset are bolded
We summarize the best models for Word2Vec and Phrase2Vec in Table 2. Overall, unigram models do the best across all three datasets both in terms of Precision and $F_{0.5}$. However, unigram Word2Vec models fail when it comes to recall often getting $2$-$13$ points less in recall as they cannot capture more complex phrases.
On the Synthetic scenario, none of the models even hit $7$-$10$ F-score, which means they are predicting most words wrong even though all sentences have a dog whistle. On Realistic and Balanced, performance significantly drops as models can only achieve $1$-$2$ F-score showing how sparser environments hurt the Word2Vec model even more.
Finally, most models do the best on the first or second level of prediction. This conservative prediction strategy decreases the usefulness of the model as it only predicts a small amount of new dog whistles. Figures displaying the choice of threshold is provided in Appendix C. Overall, unigram Word2Vec models do the best in this task but are unable to capture more complex phrases due to the limitations in the model and their conservative prediction strategy.
5.2 MLM and EPD
Scenario Model Threshold Prec DPR $F_{0.5}$
Balanced MLM Bernice 50 0.14 6.58 0.17
MLM BERTweet 50 0.23 9.47 0.28
EPD SpanBERT 50 0.00 0.00 0.00
Synthetic MLM Bernice 400 0.50 0.84 0.54
MLM BERTweet 50 2.00 0.42 1.14
EPD SpanBERT 800 1.85 3.36 2.03
Realistic MLM Bernice 6400 0.05 2.21 0.06
MLM BERTweet 400 0.50 1.47 0.58
EPD SpanBERT 200 0.58 0.74 0.61
Table 3: Results for best MLM and EPD models by $F_{0.5}$. Prec is Precision. DPR is Data Potential Recall. Best scores for across each dataset are bolded
Table 3 presents the top-performing MLM and EPD models for each scenario. Overall, MLM and EPD models perform extremely poorly across all metrics. They obtain less than $1$ point in F-score, very poor precision around $1$-$2$ points, and a modest DPR around $1$-$9$ points. For the MLM models, we attribute this lackluster performance due to limitation of the tokenizer of the models as MLM models are restricted to predicting single tokens learned from their training corpora. As for the EPD model, we believe that the filtering step using a unigram Word2Vec model leaves out complex phrases.
Looking deeper, most of the models only predict 50 words being extremely conservative with their choices which does lead to better precision, mimicking the same failure mode in Word2Vec and Phrase2Vec. Additional figures detailing the choice of threshold are provided in Appendix D.
As for model choice, EPD does the best in Synthetic and Realistic but does not work in Balanced. We hypothesize that SpanBERT is not as adapted to a domain as different as Balanced reducing its effectiveness. Overall, MLM and EPD models have lackluster performance and are beat by the simpler Word2Vec and Phrase2Vec models.
5.3 EarShot
5.3.1 PREDICT
Scenario Filtering Model Keyword Extraction Model Threshold Max $n$-grams Prec DPR $F_{0.5}$
Balanced RoBERTA R4 KeyBERT 50 1 14.81 1.65 5.70
LLaMa 13B KeyBERT 100 1 11.43 1.65 5.22
Synthetic RoBERTA R4 KeyBERT 800 1 14.86 13.45 14.55
LLaMA 8B KeyBERT 400 1 19.13 7.14 14.32
Realistic HateXplain BERT YAKE 3200 3 10.00 1.47 4.63
LLaMa 13B RAKE 200 1 9.52 1.47 4.55
Table 4: Best Results for FETCH LLM and BERT PREDICT based on $F_{0.5}$. Prec is Precision. DPR is Data Potential Recall. Best scores for across each dataset are bolded
In Table 4, we choose the best LLM and BERT filtering model for each dataset based on the $F_{0.5}$ of the system. EarShot-Predict obtains high precision of around $9$-$20$ points, beating Word2Vec, Phrase2Vec, MLM, and EPD. However, its DPR suffers obtaining around $1$-$13$ points. While it beats MLM and EPD in terms of DPR, Word2Vec and Phrase2Vec beat PREDICT on Balanced and Synthetic and tie it on Realistic. This low recall stems from the nature of keyword extraction models, which aim to identify important words rather than specifically targeting dog whistles themselves.
To achieve this high precision, the system makes the conservative choice of predicting on average less than 1000 words and only chooses $1-$grams. While its prediction abilities are still restricted, this conservative prediction threshold is an improvement over the other systems as it maintains high precision at larger thresholds, making it more competitive.
Looking towards the choices of keyword extractors, KeyBERT does the best on smaller datasets like Balanced and Synthetic and YAKE and RAKE do better on larger datasets like Realistic. For the LLM model choices, LLaMa 13B does the best across Balanced and Realistic while LLaMa 8B does better on Synthetic. As for the BERT model choices, RoBERTa R4 does best across Balanced and Synthetic while HateXplain BERT does well on Realistic. Figures detailing choices of keyword extraction models, filtering models, and thresholds can be found in Appendix E.
Comparing LLMs vs BERT for filtering, BERT-based models outperform LLMs across all three datasets by $0.1$ $0.5$ points. Even with all the extra parameters, task specific models are able to match the performance of larger general purpose LLMs for this task.
5.3.2 DIRECT
Scenario Model Prec DPR $F_{0.5}$
Balanced LLaMa 13B 2.97 13.58 3.52
LLaMa 8B 2.35 21.40 2.86
Mistral 7B 2.87 14.81 3.42
Synthetic LLaMa 13B 20.31 56.30 23.29
LLaMa 8B 14.63 46.64 16.95
Mistral 7B 18.26 58.82 21.18
Realistic LLaMa 13B 0.00 0.00 0.00
LLaMa 8B 0.94 60.29 1.17
Mistral 7B 0.69 46.32 0.86
Table 5: Results for DIRECT Pipeline. Prec is Precision. DPR is Data Potential Recall. Best scores for across each dataset are bolded
Table 5 summarizes the performance of the EarShot DIRECT system. Direct prompting of an LLM achieves high DPR but moderate precision, due to the model overestimating the number of dog whistles, inflating DPR at the cost of precision.
Diving into the models, LLaMa 13B does the best across Balanced and Synthetic while it fails on Realistic where LLaMa 8B does the best. This is due to Realistic being an order or two of magnitude larger than Balanced and Synthetic meaning LLaMa 13B overestimates the number of dog whistles per post in comparison to LLaMa 8B which is more conservative.
The DIRECT method outperforms all except PREDICT, which surpasses DIRECT on Balanced and Realistic datasets, while DIRECT excels on Synthetic in terms of $F_{0.5}$. The DIRECT method, though more powerful, fails in realistic settings. However, this result has caveats: DIRECT achieves significantly better DPR, identifies many dog whistles, is less conservative than PREDICT, and is not constrained by an $n$-gram threshold due to its generative nature. Overall, DIRECT achieves high DPR at the expense of precision, making it competitive but inferior to PREDICT.
6 Discussion
Conservative prediction maintains performance
The highest performing systems maintain a low number of predicted dog whistles. While it is useful for downstream tasks to have a small list of candidates, it hurts the ability of the model to find more dog whistles, creating a tricky trade off between recall and precision that fails to be satisfied by any of the current methods and EarShot.
EarShot beats all other methods
EarShot PREDICT and DIRECT both handily beat their competitors. They beat MLM/EPD by $1$-$18$ points in F-score and achieve higher precision and DPR across all scenarios. As for Word2Vec/Phrase2Vec, the gap is closer but still large. These performance highlight the possible usefulness of working towards creating a better version of EarShot.
Current model performance is not useful
Despite positive success with our method EarShot over other baselines, achieving around $1$-$25$ F-score across different datasets does not display a performance that could be used in the real world settings. More effort should be put into looking closer at designing a system that balances these, possibly leveraging the strengths of both EarShot-PREDICT and EarShot-DIRECT. In addition, it reflects on the limitations of LLMs as they might not be able to discover novel dog whistles without more help.
Future Directions
The findings of this study show the promise and limitations of current methodologies in identifying novel dog whistles. Most methods predict very few dog whistles to maintain high precision. While EarShot tries to improve on this case, its inability to balance precision and recall displays an area for future work. We propose multiple directions of improvement. First, we believe that hybrid EarShot-PREDICT and EarShot-DIRECT may prove fruitful, leveraging the strengths of both systems. Other improvements could include adding post processing to predicted dog whistles to increase the precision of methods, exploiting consensus methods through multiple LLMs, increasing the size of the models, prompting the model better methods through methods like chain of thought Wei et al. (2023) or ensembling external syntactic or semantic information. Ultimately, addressing this steep trade off between precision and recall will be vital for creating performant but applicable models.
7 Conclusion
This work lays the foundation for further inquiry into detecting new dog whistles using NLP. We develop a novel task to evaluate NLP systems in dog whistle discovery, rigorously test state of the art systems across multiple datasets, and propose a new system that surpasses the state of the art. We demonstrate that systems based on Word2Vec and BERT fail to achieve satisfactory performance on the task, whereas our EarShot system achieves respectable, though not perfect, performance across three distinct scenarios. We hope FETCH! will inspire future work on dog whistle discovery and creating difficult problems to push the boundary of what LLMs and NLP can do.
8 Ethical Considerations
We identify three key ethical considerations for our work. First, dog whistles are highly culturally specific. Since our study focuses on American cultural contexts, our system may misclassify words or phrases that are not considered dog whistles in other cultures, limiting its applicability and risking false positives when applied cross-culturally. Second, our system is designed to detect hidden or coded language. This ability could have the unintended consequence of exposing language that serves to protect activism or marginalized groups. Third, there is a risk that our system may exhibit bias, disproportionately classifying language from minority communities as dog whistles. Similar systems for hate speech detection have historically encountered fairness challenges Sap et al. (2019).
While we acknowledge these risks, we argue that deploying the system in real-world settings, with appropriate safeguards, can effectively mitigate potential harms. Specifically, we recommend incorporating human oversight during deployment to review model outputs, ensuring accountability, and minimizing the risks of false positives and biases. Furthermore, given the critical importance of addressing hate speech, we contend that the benefits of deploying a carefully managed system like ours outweigh the potential harms when implemented responsibly.
9 Limitations
We identify a couple key limitations of our paper. First, we work with LLMs that are expensive to run requiring fairly large GPUs to run experiments with. We try to limit ourselves to smaller but still high performance open source models in order to reduce the overhead of our system.
Second, while a majority of dog whistles are fairly unique and can be found using regex, the lack of a dedicated human labeled corpora hinders the accuracy of the benchmark as false positives could occur. We aimed to limit this problem by including the Synthetic scenario, where the dog whistles were found through different means outside of regex.
Third, data contamination may pose a concern, as LLMs trained on internet-scale datasets could have encountered the dog whistle dataset during pretraining, potentially inflating their apparent performance. More broadly, the models might have encountered these dog whistles in unlabeled contexts. Consequently, we cannot ensure that the models have not been exposed to these examples prior to evaluation.
Finally, our method has only been tried on English only datasets. However, dog whistles have not been studied in multilingual contexts as much as in English and there are less resources to study them.
Appendix A Dog Whistles Found in Each Scenario
Scenario List of Dog Whistles
Balanced #genderwoowoo,#milk,#womenwontwheesht,109,1290,1488,23/16,absent fathers,affirmative action,aiden,alarmism,alarmist,all lives matter,alt-right,america first,amerimutt,amish,anointed,attacks on or imagery of jewish financial leaders,attacks on or imagery of jewish political leaders,autoandrophile,autogynephile,baby daddies,baby mama,back the blue,bankers,banksters,barack hussein obama,bear emoji,beta,big government,big pharma,bing,biological man,biological realism,black-on-black crime,blue lives matter,blueish,boogaloo,bop,broken family,burning coal,butterfly,cabal,checkered flag emoji,china virus,clown world,clownfish,coin,colorblind,controlled media,critical race theory,cuck,cultural enrichment,cut taxes,deadbeat dad,deep state,dinosaur emojis,durden,election integrity,entitlement programs,entitlement spending,every single time,fafo,family values,federal reserve,food stamps,fren,frog emoji,gangbanger,gender critical,gender ideology,george soros,ghetto,gibsmedat,globalism,globalist,goody,google,government handout,groomers,hard-working americans,he wuz a good boy,hollywood elite,identify as,illegal aliens,illegal immigrant,illuminati,inner city,international bankers,islamic extremism,islamic extremists,islamic terrorism,islamic terrorists,islamists,israel lobby,"its okay to be white",jogger,judas,jwoke,kek,khazars,law and order,lesbophobia,"lets go brandon",lgb rights,lifelong bachelor,lizard emoji,maga,male violence,milk emoji,multiculturalism,neoliberal,new world order,nibba,npc,ok sign emoji,our guy,overpopulation,oy vey,pepe the frog,personal responsibility,pine tree emoji,political correctness,power level,prefix "super",public school,pulling strings,puppet masters,quotas,race realism,radical islam,reagan,red square emoji,references to cities with large racial minority populations being overrun by crime, drugs, rodents,religious freedom,right to work,rothschilds,safeguarding,school choice,secure the border,send me,sharia law,shlomo,silent majority,single parent,skittle,skype,snowflake,sonnenrad,soy boy,special interests,spiderweb emoji,spqr,steroids,strapping young buck,take back,thug,tough on crime,trans identified female,trans identified male,tras,triple parentheses,two lightning bolt emojis,vinland,voter fraud,war on christmas,war on drugs,welfare,white lives matter,windmill,working class,xx,yahoo,ykw,zionist,zionist occupation government
Synthetic social justice warrior, blueish, LGB rights, working class,Islamic terrorists, centipede, cuck, TRAs, globalist,Trans Identified Male, Reagan, Islamic extremists,Pepe the Frog, single parent, illegal aliens,public school, MAGA, voter fraud, soy boy, XX,gender critical, political correctness, coincidence,cut taxes, grooming, neoliberal, big pharma, Clown World,America First, government handout, New World Order,multiculturalism, references to cities with large racial minority populations being overrun by crime, drugs, rodents,beta, Islamic extremism, quotas, family values,inner city, cabal, religious freedom, cultural Marxists,steroids, NPC, deadbeat dad, 13%, fellow white people,power level, Zionist, dual citizen, 1488, Skittle,Google, race realism, Deus Vult, puppet masters, Bing,critical race theory, secure the border,Trans Identified Female, OK sign emoji, octopus,personal responsibility, actual woman, entitlement programs,baby mama, cultural enrichment, Judeo-Christian,welfare queen, gender ideology, Rothschilds, war on terror,job creators, school choice, "states rights", anointed,majority-minorit y, shekels, Barack Hussein Obama,Remove Kebab, shoah, anchor babies, ghetto,forced diversity, judicial activism, Sharia law,Federal Reserve, middle class, Hollywood elite, welfare,deep state, George Soros, illegal immigrant,affirmative action, snowflake, thug, food stamps,radical Islam, YKW, take back, law and order,the goyim know, cultural Marxism, right to work, globalism,mass immigration, bankers, genderfree, election integrity,goy, gibsmedat, Islamic terrorism, bugman,black-on-black crime, strapping young buck, biological man,frog emoji, alt-right, Kek, womyn, silent majority,Islamists, war on drugs, autogynephile, biological realism,lifelong bachelor, groomers, based, identify as,big government, entitlement spending, our guy, Yahoo,fren, "its okay to be white", Skype,institution of marriage, every single time, absent fathers,biological woman, dindu nuffin, Blue Lives Matter, Amish,special interests, triple parentheses, 41%, Pajeet,agender, hard-working Americans, forced busing,out of wedlock, broken family, property rights,safeguarding, balance the budget, international bankers,alarmist, right to know, Illuminati, male violence,overpopulation, Jwoke, All Lives Matter, autoandrophile,pulling strings, "Lets Go Brandon", baby daddies, oy vey,cosmopolitan, freedom of association, Judas,war on Christmas, hygienic, colorblind, fatherless,identitarianism, China Virus, gender socialization,controlled media, White Lives Matter, bop, early life,poisoning the well, sex is real, gangbanger,black and orange square emojis, health freedom,Zionist Occupation Government, we wuz kangz, windmill,global elite, Kalergi Plan, dindu, preemptive strike,mixed sex, economic anxiety, car salesman, genderist,brave and stunning, jogger, prefix "super", thin blue line,Sonnenrad, banksters, troon, Trilateral Commission,vaccine safety, implicit bias, central bankers,healthy tissue, erasing women, he wuz a good boy,sex-based rights, lesbian erasure, sex not gender,lesbophobia, intact, RWDS, alarmism,rootless cosmopolitan, magapede, surgical wound,burning coal, Three Percenters, gender abolitionist,three strikes laws, superpredators, SPQR,attacks on or imagery of Jewish political leaders, Dred Scott,welfare cheats, coal burner, uterus-haver, gender fandom,string puller, Boogaloo, pine tree emoji, lizard emoji,New York elite, #GenderWooWoo, New York intellectuals,protecting women and girls, merit-based immigration policy,YWNBAW, dinosaur emojis, milk emoji, "save womens sports",biological realist, cleaning up our streets, 109, COIN,Amerimutt, tough on crime, Goody, autogynephilia, Durden,coastal elite, vegan cat, FAFO, physical removal,Khazars, Aiden, pro-family, Vinland, dual loyalty,SWPL, womanface, Willie Horton, bear emoji,attacks on or imagery of Jewish financial leaders, pilpul,A Leppo, cosmopolitan elite, Holocauster, Carolus Rex,loxism, free helicopter rides, Back the Blue,adult human female, clownfish, gender abolition,abolish birthright citizenship, bix nood, war on crime,RaHoWa, cherry emoji, two lightning bolt emojis,checkered flag emoji, autoandrophilia, AFAB trans woman,butterfly, Israel Lobby, nibba, male-friendly content,open borders for Israel, Shlomo, tiny minority of men,1290, New York values
Realistic #genderwoowoo,#milk,#womenwontwheesht,109,1290,1488,23/16,absent fathers,affirmative action,aiden,alarmism,alarmist,all lives matter,alt-right,america first,amerimutt,amish,anointed,attacks on or imagery of jewish financial leaders,attacks on or imagery of jewish political leaders,autoandrophile,autogynephile,baby daddies,baby mama,back the blue,bankers,banksters,barack hussein obama,bear emoji,beta,big government,big pharma,bing,biological man,biological realism,black-on-black crime,blue lives matter,blueish,boogaloo,bop,broken family,burning coal,butterfly,cabal,checkered flag emoji,china virus,clown world,clownfish,coin,colorblind,controlled media,critical race theory,cuck,cultural enrichment,cut taxes,deadbeat dad,deep state,dinosaur emojis,durden,election integrity,entitlement programs,entitlement spending,every single time,fafo,family values,federal reserve,food stamps,fren,frog emoji,gangbanger,gender critical,gender ideology,george soros,ghetto,gibsmedat,globalism,globalist,goody,google,government handout,groomers,hard-working americans,he wuz a good boy,hollywood elite,identify as,illegal aliens,illegal immigrant,illuminati,inner city,international bankers,islamic extremism,islamic extremists,islamic terrorism,islamic terrorists,islamists,israel lobby,"its okay to be white",jogger,judas,jwoke,kek,khazars,law and order,lesbophobia,"lets go brandon",lgb rights,lifelong bachelor,lizard emoji,maga,male violence,milk emoji,multiculturalism,neoliberal,new world order,nibba,npc,ok sign emoji,our guy,overpopulation,oy vey,pepe the frog,personal responsibility,pine tree emoji,political correctness,power level,prefix "super",public school,pulling strings,puppet masters,quotas,race realism,radical islam,reagan,red square emoji,references to cities with large racial minority populations being overrun by crime, drugs, rodents,religious freedom,right to work,rothschilds,safeguarding,school choice,secure the border,send me,sharia law,shlomo,silent majority,single parent,skittle,skype,snowflake,sonnenrad,soy boy,special interests,spiderweb emoji,spqr,steroids,strapping young buck,take back,thug,tough on crime,trans identified female,trans identified male,tras,triple parentheses,two lightning bolt emojis,vinland,voter fraud,war on christmas,war on drugs,welfare,white lives matter,windmill,working class,xx,yahoo,ykw,zionist,zionist occupation government
Appendix B Prompts used in EarShot
EarShot Method Prompt
[…]
Table 7: Prompts used in EarShot
Appendix C Word2Vec/Phrase2Vec Threshold Analysis
Figure 3: Word2Vec/Phrase2Vec $F_{0.5}$ performance vs the number of words/phrases returned by the model. Plot is on log scale.
Appendix D MLM and EPD Threshold Analysis
Figure 4: MLM and EPD $F_{0.5}$ performance vs the prediction threshold. Plot is on log scale.
Appendix E EarShot PREDICT Threshold Analysis
E.1 EarShot BERT PREDICT Threshold Analysis
Figure 5: $F_{0.5}$ averaged across BERT based filtering methods displaying average performance across keyword extraction models for EarShot BERT PREDICT vs prediction threshold across all three datasets. Plot is on log scale.
Figure 6: $F_{0.5}$ averaged across keyword extraction models displaying average performance across BERT based filtering models for EarShot BERT PREDICT vs prediction threshold across all three datasets. Plot is on log scale.
E.2 EarShot LLM PREDICT Threshold Analysis
Figure 7: $F_{0.5}$ averaged across LLM based filtering methods displaying average performance across keyword extraction models for EarShot LLM PREDICT vs prediction threshold across all three datasets. Plot is on log scale.
Figure 8: $F_{0.5}$ averaged across keyword extraction models displaying average performance across LLM based filtering models for EarShot LLM PREDICT vs prediction threshold across all three datasets. Plot is on log scale.